[Hallc_running] [New Logentry] Hall C update (Dec. 10)

gaskelld at jlab.org gaskelld at jlab.org
Sun Dec 10 09:50:02 EST 2017

Logentry Text:
Progress since yesterday:

1. Centered the beam on the carbon hole target (x=-0.5, y=3.5 on IPM3H07C): https://logbooks.jlab.org/entry/3501795
2. Some time was spent checking the trigger/DAQ timing.
3. Started taking some detector checkout data (with and without sieve slit) in the SHMS and HMS.

Problems that need to be resolved:

1. Ion chamber calibration is problematic. The MCC procedure doesn't seem to converge properly - perhaps related to  EPICS readbacks.  This is making it difficult to run higher currents (https://logbooks.jlab.org/entry/3501989). SSG is aware of this problem and will look into it later today.
2. The target motion FSD will not clear when either the 4cm or 10 cm hydrogen targets are in. SSG thinks this signal is generated by the target hardware.
3. The superharp traces on 3H07A and 3H07B don't make sense (not enough peaks, show up in weird places). I thought this might be related to the fact that the big Moller quads were set to the incorrect current - even after correction, still only 2 peaks are visible. We need to look at the YAG viewer to make sure the current adjustment had the desired effect.

Right now we are down for work related to bringing Hall A up at 3 pass. The plan is that this interruption should be about 30 minutes to an hour.

The plan for the rest of the day will depend on whether the superharp issues can be resolved.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3502072
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171210/e3f784be/attachment.html>

More information about the Hallc_running mailing list